OpenStudy (anonymous):

Use long division to perform the division. (Express your answer as quotient + remainder/divisor.)

OpenStudy (anonymous):

\[ x^4 + 8x^3 − 4x^2+ x-2 \over x − 2 \]

myininaya (myininaya):

x^4/x=x^3 put x^3 on top then multiply x^3(x-2) and then put it under x^4+8x^3-4x^2+x-2 and subtract

OpenStudy (anonymous):

I will send u a file it hard explant

OpenStudy (anonymous):

\[i got x^3+10x^2+16x+33+\frac{64}{x-2}\]